How long will it take for a sum of money to double itself at 10% simple interest.
Found 3 solutions by addingup, math_tutor2020, greenestamps:
Answer by addingup(3677)

Doubling time for simple interest is 1/rate, in your case:
1/0.10 = 10 years

Answer by math_tutor2020(3835)


P = initial deposit
2P = double that deposit

r = 0.10 = decimal form of the interest rate
t = number of years

A = P*(1+r*t) represents the account balance when using simple interest

So,
A = P*(1+r*t)
2P = P*(1+0.10*t)
2 = 1+0.10*t
2-1 = 0.10t
1 = 0.10t
t = 1/0.10
t = 10


Answer: 10 years

Answer by greenestamps(13327)



With simple interest, this can be solved using common sense and basic arithmetic; no mathematical formulas are needed.

With simple interest, 10% of the original amount is added each year.
For the original amount to double, 100% has to be added to the original amount.
The number of years required to add 100% to the original amount, at 10% per year, is 100/10 = 10.

ANSWER: 10 years